Please help me. I really need the answer.
The length of the longer side of a rectangle is 7 more feet than the length of the shorter side. Find the length of the shorter side that will generate an area of at least 18 ft.

Answer by checkley71(8403)

X(X+7)>=18
X^2+7X-18>=0
(X+9)(X-2)>=0
X-2=0
X=2 ANSWER FOR THE SHORTER SIDE.
PROOF:
2(2+7)>=18
2*9>=18
18>=18